# Robust Estimation of Polyserial Correlation

### v1

@@ 5 Robust estimation of polyserial correlation
While there exist general frameworks for robust estimation with exclusively continuous variables and exclusively categorical variables, none of them are applicable to mixed variables. On the one hand, robust $M$-estimation and related methods (e.g., Huber & Ronchetti,, 2009) are intended for continuous random variables, and, broadly speaking, achieve robustness by downweighting observations with extreme values. For instance, Alfons et al., (2022) use a variation of $M$-estimation, MMMM-estimation, to robustify mediation analyses against outliers, heavy-tailedness, or skewness of the observed distribution. Such estimators are not applicable to mixed data because the ordinal variable cannot take extreme values due to being categorical, nor does it admit a direct numerical interpretation to begin with. On the other hand, the theory of robust $C$-estimation (Welz,, 2024) is designed for exclusively categorical variables. $C$-estimation downweighs categories whose empirical frequency disagrees with their corresponding theoretical frequency under a postulated model. Since one variable in the polyserial model is continuous, it does not have discrete categories, so $C$-estimators cannot be applied. However, it turns out that the fundamental idea of downweighting data points that cannot be modeled sufficiently well by a postulated model can be exploited to achieve robustness through minimum density power divergence (DPD) estimation (Basu et al.,, 1998). We explain in this section how minimum DPD estimation can be applied to the polyserial model.
@@ 5.1 Maximum likelihood revisited
The observed sample, $\{(X_{i},Y_{i})\}_{i=1}^{N}$, can be uniquely characterized by a particular empirical density function, namely
@@
for $x\in\mathbb{R},y\in\mathcal{Y}$, where the indicator function $\mathds{1}\left\{E\right\}$ takes value 1 if an event $E$ is true, and 0 otherwise. The empirical density $\widehat{f}_{N}$ only takes nonzero values when evaluated at points in the observed sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$, and is zero-valued for all other points in the uncountable product space $\mathbb{R}\times\mathcal{Y}$. As such, $\widehat{f}_{N}$ is not useful for estimating the unknown density $f_{\varepsilon,XY}$ of the unknown sampling distribution $F_{\varepsilon,XY}$ in (12), for which a kernel density estimator would be more appropriate (see e.g., Ch. 24 Van der Vaart,, 1998). Yet, $\widehat{f}_{N}$ turns out useful as a theoretical object for constructing estimators for model parameters, such as $\bm{\theta}\in\bm{\Theta}$ in the case of the polyserial model. For example, the maximum likelihood estimator $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (8) can be expressed as the minimizer of the KL divergence between the empirical density $\widehat{f}_{N}$ and model density ${p}_{XY}\left(\cdot,\cdot;\bm{\theta}\right)$, which is given by
$\begin{split}{\textnormal{KL}}\left(\widehat{f}_{N}\ \big|\big|\ {p}_{XY}\left(\cdot,\cdot;\bm{\theta}\right)\right)&=\int_{\mathbb{R}}\sum_{y\in\mathcal{Y}}\widehat{f}_{N}\left(x,y\right)\log\left(\frac{\widehat{f}_{N}\left(x,y\right)}{{p}_{XY}\left(x,y;\bm{\theta}\right)}\right)\textnormal{d}x\\ &=\frac{1}{N}\sum_{i=1}^{N}\Bigg(\log\left(\widehat{f}_{N}\left(X_{i},Y_{i}\right)\right)-\log\left({p}_{XY}\left(X_{i},Y_{i};\bm{\theta}\right)\right)\Bigg),\end{split}$ (15)
where the first equality follows by definition of the KL divergence in (13) and the second equality by the definition of $\widehat{f}_{N}$ and the convention $0\log(0)=0$. Indeed, the minimizer of ${\textnormal{KL}}\left(\widehat{f}_{N}\ \big|\big|\ {p}_{XY}\left(\cdot,\cdot;\bm{\theta}\right)\right)$ with respect to $\bm{\theta}\in\bm{\Theta}$ coincides with the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (8) since the two optimization problems are equivalent due to the former being proportional to the negative log-likelihood $\sum_{i=1}^{N}\log\big({p}_{XY}\left(X_{i},Y_{i};\bm{\theta}\right)\big)$. We refer to White, (1982) for a detailed exposition of the connection between KL divergence and ML.
@@ 5.2 Minimum power divergence
$\begin{split}{\textnormal{D}_{0}}\left(g_{1}\ \big|\big|\ g_{2}\right)&=\lim_{\alpha\downarrow 0}{\textnormal{D}_{\alpha}}\left(g_{1}\ \big|\big|\ g_{2}\right)\\ &=\int\int g_{1}(s,t)\log\left(\frac{g_{1}(s,t)}{g_{2}(s,t)}\right)\textnormal{d}t\textnormal{d}s\\ &={\textnormal{KL}}\left(g_{1}\ \big|\big|\ g_{2}\right),\end{split}$ (16)
@@ 5.3 Proposed robust estimator
which is satisfied for $\bm{\theta}=\widehat{\bm{\theta}}_{N}$, and where
$\bm{c}_{\alpha}(\bm{\theta})=\int_{\mathbb{R}}\sum_{y\in\mathcal{Y}}{p}_{XY}^{1+\alpha}\left(x,y;\bm{\theta}\right)\bm{u}_{\bm{\theta}}\left(x,y\right)\textnormal{d}x$
is a correction factor independent of observed data. Observe that for $\alpha=0$, the estimating equation (19) equals that of the MLE, [^3] By elementary likelihood theory, $\bm{c}_{0}(\bm{\theta})=\int_{\mathbb{R}}\sum_{y\in\mathcal{Y}}{p}_{XY}\left(x,y;\bm{\theta}\right)\bm{u}_{\bm{\theta}}\left(x,y\right)\textnormal{d}x=\mathbb{E}_{P_{XY}}\left[\bm{u}_{\bm{\theta}}\left(X,Y\right)\right]=\bm{0}$ for all $\bm{\theta}\in\bm{\Theta}$. in which all observations in the sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$ are weighted equally. Conversely, when $\alpha>0$, the term
@@
$\widetilde{w}_{i,\alpha}(\bm{\theta})={p}_{XY}^{\alpha}\left(X_{i},Y_{i};\bm{\theta}\right).$
A weight $\widetilde{w}_{i,\alpha}(\bm{\theta})$ is bounded from below by 0 and takes values close to 0 for observations that the polyserial model cannot fit well, such as contaminated data points. As such, “observations that are widely discrepant to the [polyserial] model will get nearly zero weights” (Basu et al.,, 1998, pp. 551–52). Larger choices of $\alpha>0$ lead to more stringent downweighting. Conversely, if contamination is absent $(\varepsilon=0)$, it can be shown that the robust estimator $\widehat{\bm{\theta}}_{N}$ and the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (8) are asymptotically equal to one another for all choices $\alpha\geq 0$. This property is implied by the property of Fisher consistency that minimum DPD estimators posses (implied by Theorem 1 in Basu et al.,, 1998, see p. 551 therein).
By construction, for $\alpha>0$, the robust estimator’s raw weights $\widetilde{w}_{i,\alpha}(\bm{\theta})$ are bounded from above by some data-independent finite value $M_{\alpha}(\bm{\theta})$ only depending on $\alpha$ and $\bm{\theta}$, which is defined by $M_{\alpha}(\bm{\theta})=\sup\{{p}_{XY}^{\alpha}\left(x,y;\bm{\theta}\right):x\in\mathbb{R},y\in\mathcal{Y}\}$. We believe that it appealing for interpretation if the weights of perfectly fitting observations receive a value of exactly 1. We therefore rescale the raw weights by their upper bound $M_{\alpha}(\bm{\theta})$ to obtain
$w_{i,\alpha}(\bm{\theta})=\widetilde{w}_{i,\alpha}(\bm{\theta})\big/M_{\alpha}(\bm{\theta}),\qquad i=1,\dots,N,$ (20)
and henceforth refer to the rescaled weights $w_{i,\alpha}(\bm{\theta})$ simply as weights. The rescaled weights are contained in the interval $[0,1]$ and take values close to 0 for poorly fitting observations and value 1 for perfectly fitting observations. This rescaling is without loss of generality because the estimating equation (19) can equivalently be expressed as
@@
Appendix B describes an algorithm for computing the upper bound $M_{\alpha}(\bm{\theta})$.
@@ 6.1 Estimand
On the other hand, if contamination is present $(\varepsilon>0)$, then the estimand $\bm{\theta}_{0}$ will generally differ from the true $\bm{\theta}_{*}$. How much they differ depends on the contamination fraction $\varepsilon$ and contamination type $H_{X\eta}$ in the contaminated distribution (11), as well as the choice of tuning constant $\alpha\geq 0$. Roughly speaking, holding the unknown contaminated distribution $F_{\varepsilon,X\eta}$ constant, the larger $\alpha$, the closer $\bm{\theta}_{0}$ is to the true $\bm{\theta}_{*}$. In other words, if the polyserial model is misspecified, larger values of $\alpha$ asymptotically lead to less bias for the true $\bm{\theta}_{*}$.
@@ 6.3 Efficiency
where the operator $(\cdot)_{1,1}$ picks out the top left element of a matrix.
@@
Table 1 lists the relative efficiencies of various choices of the tuning constant $\alpha$ for an ordinal $Y$ with $r=5$ response options and true parameter vector $\bm{\theta}_{*}=\left(\rho_{*},\mu_{*},\sigma_{*}^{2},\tau_{*,1},\tau_{*,2},\tau_{*,3},\tau_{*,4}\right)^{\top}=\left(0.5,0,1,-1.5,-0.5,0.5,1.5\right)^{\top}$. Up to about $\alpha=0.25$, the relative efficiency of the robust estimator stays above 90%, and then decreases roughly linearly to slightly below 50% for $\alpha=1$. Choices of $\alpha>1$, while yielding even more robustness, would result in a comparatively poor relative efficiency of considerably less than 50% and are therefore not considered in this paper. The efficiency results for other choices of the true parameter vector $\bm{\theta}_{*}$ are very similar and are therefore deferred to Appendix C (Figure C.1). Our software implementation, discussed in more detail in Section 6.4, provides functionality to compute efficiency at arbitrary user-specified parameter vectors and tuning constants $\alpha$.
A loss of efficiency is a common property of many robust estimators involving continuous random variables, and, as demonstrated in this section, our proposed estimator is no exception. [^5] In contrast, for exclusively categorical variables, it is possible to construct robust estimators that are fully efficient. See Welz, (2024) for general models of categorical data, and Welz et al., 2025b for the special case of the polychoric correlation model. As such, our estimator is based on a compromise between efficiency and robustness. Borrowing an insurance metaphor from Anscombe, (1960), we “sacrifice some efficiency at the model, in order to insure against accidents caused by deviations from the model” (Huber & Ronchetti,, 2009, p. 5). However, it turns out that the efficiency loss is relatively minor for reasonably small positive tuning constants like $\alpha=0.1$ (Table 1), which, as the simulation experiments in Section 7 will reveal, suffices to gain substantial robustness over ML estimation.
@@ 6.4 Implementation
We provide am open-source implementation of our robust estimator as part of the package robcat (for “ROBust CATegorical data analysis”; Welz et al., 2025a, ) for the statistical programming environment R (R Core Team,, 2024). The package is freely available from CRAN (the Comprehensive R Archive Network) at https://CRAN.R-project.org/package=robcat. We used this package for obtaining all numerical results in this paper.
In principle, every suitable method for numerical optimization can be used to minimize the robust estimator’s minimization problem in (18). In our experience, unconstrained optimization via the BFGS algorithm (e.g., Nocedal & Wright,, 2006, Section 6.1) works well. Yet, additional stability could be gained from making explicit the constraints in the parameter space $\bm{\Theta}$ in (9), for which standard algorithms for constrained optimization can be used, such as the simplex method of Nelder & Mead, (1965). In our implementation, we adopt a similar default behavior as the implementation of robust polychoric correlation estimation (Welz et al., 2025b, ) in package robcat. Specifically, the implementation first tries unconstrained optimization with the BFGS algorithm. If instability or numerical nonconvergence are encountered or a constraint is violated, it instead uses the algorithm of Nelder & Mead, (1965) for constrained optimization. Nevertheless, numerous other optimization algorithms are supported, and users can freely choose their preferred option.

### v2

@@ 5 Robust estimation of polyserial correlation
While there exist general frameworks for robust estimation with exclusively continuous variables and exclusively categorical variables, none of them are applicable to mixed variables. On the one hand, robust $M$-estimation as proposed by Huber, (1964) and related methods (e.g., Huber & Ronchetti,, 2009) are intended for continuous random variables, and, broadly speaking, achieve robustness by downweighting observations with extreme values. For instance, Alfons et al., (2022) use a variation of $M$-estimation, MMMM-estimation, to robustify mediation analyses against outliers, heavy-tailedness, or skewness of the observed distribution. Such estimators are not applicable to mixed data because the ordinal variable cannot take extreme values due to being categorical, nor does it admit a direct numerical interpretation to begin with. On the other hand, the theory of robust $C$-estimation (Welz,, 2024) is designed for exclusively categorical variables. $C$-estimation downweighs categories whose empirical frequency disagrees with their corresponding theoretical frequency under a postulated model. Since one variable in the polyserial model is continuous, it does not have discrete categories, so $C$-estimators cannot be applied. However, it turns out that the fundamental idea of downweighting data points that cannot be modeled sufficiently well by a postulated model can be exploited to achieve robustness through minimum density power divergence (DPD) estimation (Basu et al.,, 1998). We explain in this section how minimum DPD estimation can be applied to the polyserial model.
@@ 5.1 Maximum likelihood revisited
The observed sample can be uniquely characterized by a particular empirical density function,
@@
for $x\in\mathbb{R},y\in\mathcal{Y}$, where the indicator function $\mathds{1}\left\{E\right\}$ takes value 1 if an event $E$ is true, and 0 otherwise. As such, the empirical density $\widehat{f}_{N}$ only takes nonzero values when evaluated at points in the observed sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$. The maximum likelihood estimator $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (5) can be expressed as the minimizer of the KL divergence between the empirical density $\widehat{f}_{N}$ and model density ${p}_{XY}\left(\cdot,\cdot;\bm{\theta}\right)$, which is given by
${\textnormal{KL}}\left(\widehat{f}_{N}\ \big|\big|\ {p}_{XY}\left(\cdot,\cdot;\bm{\theta}\right)\right)=\frac{1}{N}\sum_{i=1}^{N}\Bigg(\log\left(\widehat{f}_{N}\left(X_{i},Y_{i}\right)\right)-\log\left({p}_{XY}\left(X_{i},Y_{i};\bm{\theta}\right)\right)\Bigg),$ (12)
and follows by definition of $\widehat{f}_{N}$ and the KL divergence in (10) as well as the convention $0\log(0)=0$. [^5] Alternatively, the ML estimator (and also the proposed robust estimator) can be equivalently defined using statistical divergence functionals, as done in Basu et al., (1998). We refer to White, (1982) for a detailed exposition of the connection between KL divergence and ML.
@@ 5.2 Minimum power divergence
${\textnormal{D}_{0}}\left(g_{1}\ \big|\big|\ g_{2}\right)=\lim_{\alpha\downarrow 0}{\textnormal{D}_{\alpha}}\left(g_{1}\ \big|\big|\ g_{2}\right)=\int\int g_{1}(s,t)\log\left(\frac{g_{1}(s,t)}{g_{2}(s,t)}\right)\textnormal{d}t\textnormal{d}s={\textnormal{KL}}\left(g_{1}\ \big|\big|\ g_{2}\right),$ (13)
@@ 5.3 Proposed robust estimator
which is satisfied for $\bm{\theta}=\widehat{\bm{\theta}}_{N}$ and where $\bm{c}_{\alpha}(\bm{\theta})=\int_{\mathbb{R}}\sum_{y\in\mathcal{Y}}{p}_{XY}^{1+\alpha}\left(x,y;\bm{\theta}\right)\bm{u}_{\bm{\theta}}\left(x,y\right)\textnormal{d}x$ is a correction factor independent of observed data. Observe that for $\alpha=0$, the estimating equation (16) equals that of the MLE, [^6] By elementary likelihood theory, $\bm{c}_{0}(\bm{\theta})=\int_{\mathbb{R}}\sum_{y\in\mathcal{Y}}{p}_{XY}\left(x,y;\bm{\theta}\right)\bm{u}_{\bm{\theta}}\left(x,y\right)\textnormal{d}x=\mathbb{E}_{P_{XY}}\left[\bm{u}_{\bm{\theta}}\left(X,Y\right)\right]=\bm{0}$ for all $\bm{\theta}\in\bm{\Theta}$. in which all observations in the sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$ are weighted equally. Conversely, when $\alpha>0$, the term
@@
$\widetilde{w}_{i,\alpha}(\bm{\theta})={p}_{XY}^{\alpha}\left(X_{i},Y_{i};\bm{\theta}\right).$ (17)
A weight $\widetilde{w}_{i,\alpha}(\bm{\theta})$ is bounded from below by 0 and takes values close to 0 for observations that the polyserial model cannot fit well, such as contaminated data points. As such, “observations that are widely discrepant to the [polyserial] model will get nearly zero weights” (Basu et al.,, 1998, pp. 551–52). Larger choices of $\alpha>0$ lead to more stringent downweighting. For instance, a contaminated observation with an ordinal response $Y_{i}$ that strongly disagrees with the latent threshold structure implied by the model density $p_{XY}$ would receive a weight close to 0.
If contamination is absent $(\varepsilon=0)$, it can be shown that the robust estimator $\widehat{\bm{\theta}}_{N}$ and the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (5) are asymptotically equal to one another for all choices $\alpha\geq 0$. This property is implied by the property of Fisher consistency that minimum DPD estimators posses (implied by Theorem 1 in Basu et al.,, 1998, see p. 551 therein).
Being expressible as a root of a summation over the sample where each summand is a certain function evaluated at a single observation (Eq. 16), minimum DPD estimators are a special case of $M$-estimation (see Basu et al.,, 1998, Section 3.1). However, unlike classic $M$-estimation as proposed in Huber, (1964), minimum DPD estimation does not require the data to admit a numerical interpretation due to operationalizing “outlyingness” through discrepant model fit rather than extreme values. Consequently, it can be directly applied to models for mixed data, such as the polyserial model.
5.4 Rescaling of weights
By construction, for a given tuning constant $\alpha>0$, the robust estimator’s raw weights $\widetilde{w}_{i,\alpha}(\bm{\theta})$ in (17) are bounded from below by 0 and from above by some data-independent finite value $M_{\alpha}(\bm{\theta})$ only depending on $\alpha$ and $\bm{\theta}$, which is defined by $M_{\alpha}(\bm{\theta})=\sup\{{p}_{XY}^{\alpha}\left(x,y;\bm{\theta}\right):x\in\mathbb{R},y\in\mathcal{Y}\}$. The fact that the upper bound $M_{\alpha}(\bm{\theta})$ varies for different values of $\bm{\theta}$ and $\alpha$ makes it difficult to compare raw weights across different estimates and specifications. Since the raw weights are intended as a tool for detecting contaminated observations, it would be beneficial to have such a comparability. In the following, we describe a simple novel procedure to rescale the weights to always be bounded from above by 1. Rescale the raw weights by their upper bound $M_{\alpha}(\bm{\theta})$ to obtain
$w_{i,\alpha}(\bm{\theta})=\widetilde{w}_{i,\alpha}(\bm{\theta})\big/M_{\alpha}(\bm{\theta}),\qquad i=1,\dots,N.$ (18)
We henceforth refer to the rescaled weights $w_{i,\alpha}(\bm{\theta})$ simply as weights. The rescaled weights are contained in the interval $[0,1]$ and take values close to 0 for poorly fitting observations and value 1 for perfectly fitting observations. This rescaling is without loss of generality because the estimating equation (16) can equivalently be expressed as
@@ 5.4 Rescaling of weights
Appendix B describes an algorithm for computing the upper bound $M_{\alpha}(\bm{\theta})$. The usefulness of the rescaled weights as tool for identifying contaminated data points will be demonstrated in an empirical application in Section 8, where they successfully detect observations discrepant from the model.
@@ 6.1 Estimand
On the other hand, if contamination is present $(\varepsilon>0)$, then the estimand $\bm{\theta}_{0}$ will generally differ from the true $\bm{\theta}_{*}$. How much they differ depends on the contamination fraction $\varepsilon$ and contamination type $H_{X\eta}$ in the contaminated distribution (8), as well as the choice of tuning constant $\alpha\geq 0$. Roughly speaking, holding the unknown contaminated distribution $F_{\varepsilon,X\eta}$ constant, the larger $\alpha$, the closer $\bm{\theta}_{0}$ tends to be to the true $\bm{\theta}_{*}$. [^8] We stress that this sentence is a heuristic rather than a formal result. Since no assumption is made on the contamination distribution $H_{X\eta}$, there may exist contaminated distributions $F_{\varepsilon,X\eta}$ for which higher values of $\alpha$ do not uniformly reduce bias. We refer to Hampel et al., (1986, Chapter 8.2d) for a discussion on the fundamental difficulty of deriving bias-reduction results of robust procedures in contamination models such as (8). In other words, if the polyserial model is misspecified, larger values of $\alpha$ asymptotically tend to lead to less bias for the true $\bm{\theta}_{*}$.
@@ 6.3 Efficiency
where the operator $(\cdot)_{1,1}$ picks out the top left element of a matrix. The efficiency loss of minimum DPD estimators can also be expressed as a function of $\alpha$ for a given model (see Basu et al.,, 1998, Section 4.2).
@@
Table 1 lists the relative efficiencies of various choices of the tuning constant $\alpha$ for an ordinal $Y$ with $r=5$ response options and true parameter vector $\bm{\theta}_{*}=\left(\rho_{*},\mu_{*},\sigma_{*}^{2},\tau_{*,1},\tau_{*,2},\tau_{*,3},\tau_{*,4}\right)^{\top}=\left(0.5,0,1,-1.5,-0.5,0.5,1.5\right)^{\top}$. Up to about $\alpha=0.25$, the relative efficiency of the robust estimator stays above 90%, and then decreases roughly linearly to slightly below 50% for $\alpha=1$. Choices of $\alpha>1$, while yielding even more robustness, would result in a comparatively poor relative efficiency of considerably less than 50% and are therefore not considered in this paper. The efficiency results for other choices of the true parameter vector $\bm{\theta}_{*}$ are very similar and are therefore deferred to Appendix C (Figure C.1). Our software implementation, discussed in more detail in Section 6.4, provides functionality to compute the relative efficiency at arbitrary user-specified parameter vectors and tuning constants $\alpha$.
A loss of efficiency is a common property of many robust estimators involving continuous random variables, and, as demonstrated in this section, our proposed estimator is no exception. [^9] In contrast, for exclusively categorical variables, it is possible to construct robust estimators that are fully efficient. See Welz, (2024) for general models of categorical data, and Welz et al., (2026) for the special case of the polychoric correlation model. As such, our estimator is based on a compromise between efficiency and robustness. Borrowing an insurance metaphor from Anscombe, (1960), we “sacrifice some efficiency at the model, in order to insure against accidents caused by deviations from the model” (Huber & Ronchetti,, 2009, p. 5). However, it turns out that the efficiency loss is relatively minor for reasonably small positive tuning constants like $\alpha=0.1$ (Table 1), which, as the simulation experiments in Section 7 will reveal, suffices to gain substantial robustness over ML estimation.
@@ 6.4 Implementation
We provide an open-source implementation of our robust estimator as part of the package robcat (for “ROBust CATegorical data analysis”; Welz et al.,, 2025) for the statistical programming environment R (R Core Team,, 2024). The package is freely available from CRAN (the Comprehensive R Archive Network) at https://CRAN.R-project.org/package=robcat. We used this package for obtaining all numerical results in this paper.
In principle, every suitable method for numerical optimization can be used to minimize the robust estimator’s minimization problem in (15). In our experience, unconstrained optimization via the BFGS algorithm (e.g., Nocedal & Wright,, 2006, Section 6.1) works well. Yet, additional stability could be gained from making explicit the constraints in the parameter space $\bm{\Theta}$ in (6), for which standard algorithms for constrained optimization can be used, such as the simplex method of Nelder & Mead, (1965). In our implementation, we adopt a similar default behavior as the implementation of robust polychoric correlation estimation (Welz et al.,, 2026) in package robcat. Specifically, the implementation first tries unconstrained optimization with the BFGS algorithm. If instability or numerical nonconvergence are encountered or a constraint is violated, it instead uses the algorithm of Nelder & Mead, (1965) for constrained optimization. [^10] In our experience, the higher the contamination fraction, the more likely it is that unconstrained optimization fails, so constrained algorithms must be used instead. As an example, we refer to the convergence statistics for the simulation in Section 7, reported in Appendix C. Nevertheless, numerous other optimization algorithms are supported, and users can freely choose their preferred option.